H jumps upward off the end of a 1.0m high diving board into the pool. At the same time, K steps off the adjacent 3.0m high board. What is H's initial speed if they both hit the water at the same time?
Answer #1

Both hit at the same time

speed of K when it hits

v = sqrt(2gh) = sqrt(2*9.8*3) = 7.668 m/s

time taken by k

v = u + gt

7.668 = 0 + 9.8*t

t = 0.78 s

time taken by K = 0.78s

so time taken by H will also be 0.78 s

H = ut + 0.5gt^2

1 = u*0.78 + 0.5*-9.8*0.78^2

u = 5.12 m/s

so H's initial speed is 5.12 m/s